22/23 Suppose the marginal cost function for a month for a certain product is:               MC...

22/23

Suppose the marginal cost function for a month for a certain product is:

              MC = 3X + 50, where X = # of units; cost in $. If the fixed costs = $10,000          per month. Find the total cost function for the month.

The rate of change of the cost (marginal cost) for a product is:

            MC = 3(2X+25)^1/2;    X=#units; Cost in $

           If the fixed costs per month = $11,125, what would be the total cost of producing 300 items per month.
Answer #1

@ Given: MC = 3 (22+28) Jde - J5 (224293 de = 3 (22423 % te - 32 I c = 2 (287253912 + 11,125 at x = 300, c = 2 ( 2830 +25) %